If the sides of a square are lengthened by 6cm, the area becomes 144cm^2. Find the length of a side of the original square.
Please explain answer. Thanks!

Answer by ankor@dixie-net.com(22740)

If the sides of a square are lengthened by 6cm, the area becomes 144cm^2.
Find the length of a side of the original square.
:
Let x = side of the original square
then
(x+6) = lengthened side of the square
:
The area of the lengthened square is given as 144, therefore
(x+6)^2 = 144
Find the square root of both sides
x + 6 = 12
x = 12 - 6
x = 6 cm is the side of the original square